log 2 (x 2 -3x-2)>= 4 then x lies in the interval a) [6,∞) b) (-∞,-1) c) [6,10] d) [-1,-6]

log2(x2-3x-2)>= 4 then x lies in the interval
a) [6,∞)
b) (-∞,-1)
c) [6,10]
d) [-1,-6]

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
X² - 3x -2 greater than 0
And x² +3x-2 greater than or equal to 16
 
Take intersection of these two solution. You will get your answer
Aditya Gupta
4 years ago
log2(x2-3x-2)>= 4 
raise both sides to power of 2.
2^log2(x2-3x-2)>= 2^4 
x^2 – 3x – 2 >= 16
or (x – 6)(x+3)>= 0
or x belongs to ( – inf, – 3]U[6, inf).
option a and c are correct. seems like a multi correct problem. anyway the complete soln is x belongs to ( – inf, – 3]U[6, inf).
KINDLY APPROVE :))